# A Controllable Examination for Long-Context Language Models

### v1

@@ Abstract
Existing frameworks for evaluating long-context language models (LCLM) can be broadly categorized into real-world and synthetic tasks. Despite their utility, both approaches are accompanied by certain intrinsic limitations. For example, real-world tasks are too complex to interpret or characterize and are susceptible to data contamination. In contrast, synthetic tasks often adopt the needle-in-the-haystack (NIAH) format, wherein a lack of coherence between the “needle” and the “haystack” compromises their validity as proxies for realistic applications. In response to these challenges, we posit that an ideal long-context evaluation framework should be characterized by three essential features: seamless context, controllable setting, and sound evaluation. This study introduces LongBioBench, a novel benchmark that utilizes artificially generated biographies as a controlled environment for assessing LCLMs across dimensions of understanding, reasoning, and trustworthiness. Our experimental evaluation, which includes 18 LCLMs in total, demonstrates that most models still exhibit deficiencies in semantic understanding and elementary reasoning over retrieved results and are less trustworthy as context length increases. Our further analysis indicates some design choices employed by existing synthetic benchmarks, such as contextual non-coherence, numerical needles, and the absence of distractors, rendering them vulnerable to test the model’s long-context capabilities. Moreover, we also reveal that long-context continual pretraining primarily adjusts RoPE embedding to accommodate extended context lengths, which in turn yields only marginal improvements in the model’s true capabilities. To sum up, compared to previous synthetic benchmarks, LongBioBench achieves a better trade-off between mirroring authentic language tasks and maintaining controllability, and is highly interpretable and configurable.  [^1] The code and data are available at https://github.com/Thomasyyj/LongBio-Benchmark.
@@ 1 Introduction
Existing long-context evaluation benchmarks are primarily of two types: natural and synthetic. Regarding benchmarks with natural datapoints,  Karpinska et al. [11] collects different novels read by annotators and asks them to annotate true/false of the narrative phenomenon.  Li et al. [17] collects the latest documents post year 2022 and crafts questions for them. Though they provide authentic language samples, they are expensive to collect and annotate and susceptible to data contamination. Moreover, their inherent complexity makes them hard to characterize for controlled studies. Thus, the model’s performance on such benchmarks often fails to shed sufficient light on the model’s underlying bottlenecks. In other words, while the data is genuine, it does little to pinpoint precisely how and why a model might struggle, offering limited insights into how its long-context capabilities can be improved. Moreover, since their questions are crafted by human annotators, the task is frozen after annotation and thus unextensible to new evaluation scenarios.
On the contrary, synthetic benchmarks are highly controllable. One can isolate specific aspects they want to test with carefully designed synthetic tasks. Benchmarks like RULER [14] allow practitioners to systematically adjust variables and examine their targeted hypotheses about the model’s potential behaviors. However, existing synthetic benchmarks mostly follow the Needle-In-The-Haystack (NIAH) [18] format, where the context is usually non-coherent because the needle (the information to be recalled) is semantically unrelated to the continuous context. We show in Sec. 4.2 that this makes the benchmark less challenging when the tasks become more difficult, as it potentially presents shortcuts for the model to retrieve the target information. Meanwhile, the complexity of the needle itself is another limitation. As we will show in Sec. 4.1, the numerical needle, which is employed by NIAH [18] and RULER [14] benchmarks, is much easier to retrieve than other types of information. This implicit bias makes them worse proxies of real-world applications, hindering the development of stronger models.
@@
Inspired by Allen-Zhu and Li [19], this paper proposes LongBioBench, employing fictional biographies as a playground to examine existing long-context language models. In our framework, a single configurable biography (the “needle”) is seamlessly embedded within a set of other configurable biographies (the “haystack”), enabling diverse and scalable task design grounded in each bio’s rich factual content. Notably, we show in Fig. 1 that the evaluation scores on our purely synthetic benchmark exhibit a stronger correlation (0.853) with the scores of HELMET [12], which employs real-world tasks, than with other existing purely synthetic dataset RULER [14] (0.559).
@@ 2.2 Data Construction
We use the simplest version, referred to as Standard version, as an example to show how a single data point is constructed. An illustration of our proposed context is shown on the left of Fig. 2, and the detailed construction progress is explained in App. A. Specifically, given a task, we first generate needle and haystack biographies using a biography generator. The needle biography is inserted into the haystack to form the context. The corresponding questions and answers are generated alongside the needle biographies, yielding a complete data point. For the biography generator, it typically samples six attributes from predefined attribute pools and fills them into human-written templates to produce coherent biographies for each individual. We also provide flexible configuration options at each step of the construction process to ensure controllability. For instance, we can adjust the distractor density when generating the haystack bios or define the number of required needles when constructing the needle bios. We report the statistics of the dataset in App. B.
The proposed benchmark closely matches the desiderata aforementioned. For seamless context, our generated biographies are ensured to be naturally fluent and maintain the coherence of the entire context. Regarding controllability, the benchmark is highly modular and configurable, allowing us to do isolated ablation studies to probe the model’s behavior. The benchmark is also extensible, as the richly factual nature of the context enables a wide range of downstream tasks, depending on how the embedded knowledge is manipulated (further discussed in the next section). Finally, we ensure sound evaluation because the final answer in our bench can be easily verified. We also prevent contamination because all biographies are artificial and can be generated on-the-fly, thus avoiding reliance on any parametric knowledge memorized by the model.
@@ 2.3 Task Description
This subsection will explain how we extend to have all the tasks in LongBioBench. The tasks are split into three categories: understanding, reasoning, and trustworthiness, representing the core capabilities required to solve the tasks. An overview of how each task is extended is shown on the right of Fig. 2 and the detailed description for each task is in App. C .
@@
Long Context Understanding
Long Context Reasoning
Trustworthiness
3 Main Evaluation Results
3.1 Evaluation setup
We evaluate 15 open-source LCLMs supporting more than 128k context lengths, such as Llama [21], Phi [22], Qwen2.5 [23], Mistral [24]. Details are summerized in App. E. We also include three closed-source models in the GPT family [25]. Each model is evaluated at input lengths: $L\in\{2K,8K,16K,32K,64K,128K\}$ where L is the number of tokens counted with the tokenizer of each model. We use zero-shot prompts in all understanding and IDK tasks, and use 2-shot prompting for all reasoning and citation tasks to ensure that the model follows the answer format. We list the prompts for all tasks in App. G. During initial studies, we observed that the model’s performance on our proposed tasks nearly converges  [^2] The accuracy fluctuates within 1% for 32 data points at around 800 data points, so each test set contains 800 data points. We use the vLLM [26] framework for inference with 8$\times$×H800 GPUs. We use greedy decoding for all models. Regarding evaluation metrics, we use accuracy with exact match for all understanding and reasoning tasks (all-or-nothing accuracy is applied in the multiple retrieval case). We measure the accuracy of the citations only in the citation task. For the IDK task, the metric is the proportion of questions where the model answers correctly when information is present and refuses to answer when it is absent.
3.2 Validating all proposed tasks
3.3 Challenges for current long context modeling
This section summarizes six key results observed when evaluating LCLMs on LongBioBench. We first analyze the overall performance across all LCLMs from Fig. 3 (R1 and R2) and look deeper into the performance of different tasks in Fig. 4 (R3-R6). The full results are shown in App. F.
R3: Context length is still the main bottleneck.
R4: Poor correlation between calculation and other reasoning tasks
R5: LCLMs struggle on constrained planning problem
We construct twodiff as an example of a hard task for LCLMs. The answer to this task is not determined, and all bios in the context could serve as the needle. As shown in Fig. 4, even models with strong multihop and arithmetic reasoning abilities—such as Qwen-2.5-14B-1M—struggle with constrained retrieval, failing to perform well even at a 2k context length. Besides, none of the models perform over 30% accuracy at 128k context level, which could be easier with a bigger selection pool. This aligns with the findings from [13] that LCLMs only use a small part of context when doing the reasoning task. This highlights a fundamental limitation: current LCLMs remain far from being able to reason effectively over long contexts.
4 Analysis
4.2 Coherent context is important
4.3 Performance trend during long-context continual pre-training
Setting
Performance saturates at the early stage.
Long context pretraining boosts retrieving, but does not help reasoning
Comparing the middle figure with the left one, we observe that performance improves consistently across all retrieval tasks, whereas the reasoning task shows only a slight improvement, with accuracy remaining extremely low at around 10% (note that the random guess accuracy for the ranking task is 50% since it involves ranking two individuals). This indicates that pretraining on longer contexts primarily enhances retrieval capabilities but not reasoning abilities. Interestingly, the calculation task follows a performance trajectory similar to all retrieval tasks and already achieves high accuracy before long-context pretraining. This suggests that Qwen2.5 already possesses the capability to perform calculations over relatively long contexts, and that the main bottleneck on this task lies in retrieval rather than reasoning. Therefore, we categorize the calculation task alongside the retrieval tasks.
The model becomes less trustworthy as the training proceeds
4.4 Distractor density is another bottleneck for long context tasks
To further investigate the factors influencing the performance of LCLMs, we conduct a stress test on Qwen2.5-Instruct-7B-1M by controlling the position of the answer information and the density of distractors as the variables while keeping the context length and task fixed. Detailed analysis is shown in App. D and we summarize the results as follows: (1) We observe a strong negative correlation between distractor density and model performance, suggesting that beyond context length, higher distractor density is a key factor contributing to the difficulty LCLMs face with long-context tasks. (2) We observe the lost-in-the-middle [28] but it is less evident on relatively easier tasks.
5 Related Work
With the growing interest in long-context language modeling, various benchmarks have emerged. Some focus on real-world tasks across diverse domains, such as document understanding [11, 29, 27], safety [30], and medical question answering [31]. Others try to collect a collection of real-world tasks aimed at a comprehensive evaluation [10, 32, 12]. However, these benchmarks are often costly to construct and suffer from limited interpretability, as it is challenging to control task complexity systematically. On the other hand, many synthetic tasks are proposed due to their low construction cost and high flexibility [14, 33, 34, 13, 15, 35]. Among them, Needle-in-a-Haystack (NIAH) [18] is the most popular setting, where a “needle” is inserted at a long essay (i.e., the haystack) and the model is tasked with retrieving it. There are also variants of NIAH such as RULER [14]. Nonetheless, several studies have raised concerns about the limitations of this approach [13, 16]. In this work, we argue that the semantic irrelevance between the needle and the surrounding context may allow models to exploit cues or shortcuts, ultimately reducing the challenge and bringing bias into evaluations.
6 Conclusion and Limitations
Appendix A Dataset Construction Details
Attribute Sampling
Bio construction
Context Synthesis
Appendix B Data Statistics
Task/Length 2 8 16 32 64 128
standard 14.93 73.77 152.19 308.83 622.18 1248.75
paraphrase 12.88 62.36 128.01 263.86 528.61 1060.80
pronoun 15.07 75.21 156.65 316.81 636.10 1276.81
multi_standard 12.05 70.87 149.28 305.35 617.94 1246.20
calculation 12.84 76.34 161.12 330.70 668.71 1348.31
rank 12.86 75.77 160.62 329.90 667.42 1345.70
multihop 12.06 70.88 149.27 305.82 619.39 1246.22
twodiff 12.85 76.31 161.42 330.84 670.33 1347.75
Task/Length 2 8 16 32 64 128
paraphrase $126.85_{12.68}$ $125.16_{12.15}$ $125.59_{11.60}$ $123.48_{11.45}$ $124.14_{11.23}$ $124.14_{11.21}$
Appendix C Task Description
The tasks are split into three categories: understanding, reasoning, and trustworthiness, representing the core capabilities required to solve the tasks. An overview of the benchmark is presented in Table 4.
Table 4: Task Overview for the LongBioBench

### v2

@@ Abstract
Existing frameworks for evaluating long-context language models (LCLM) can be broadly categorized into real-world applications (e.g, document summarization) and synthetic tasks (e.g, needle-in-a-haystack). Despite their utility, both approaches are accompanied by certain intrinsic limitations. Real-world tasks often involve complexity that makes interpretation challenging and suffer from data contamination, whereas synthetic tasks frequently lack meaningful coherence between the target information ("needle") and its surrounding context ("haystack"), undermining their validity as proxies for realistic applications. In response to these challenges, we posit that an ideal long-context evaluation framework should be characterized by three essential features: 1) seamless context: coherent contextual integration between target information and its surrounding context; 2) controllable setting: an extensible task setup that enables controlled studies—for example, incorporating additional required abilities such as numerical reasoning; and 3) sound evaluation: avoiding LLM-as-Judge and conduct exact-match to ensure deterministic and reproducible evaluation results. This study introduces LongBioBench, a benchmark that utilizes artificially generated biographies as a controlled environment for assessing LCLMs across dimensions of understanding, reasoning, and trustworthiness. Our experimental evaluation, which includes 18 LCLMs in total, demonstrates that most models still exhibit deficiencies in semantic understanding and elementary reasoning over retrieved results and are less trustworthy as context length increases. Our further analysis indicates some design choices employed by existing synthetic benchmarks, such as contextual non-coherence, numerical needles, and the absence of distractors, rendering them vulnerable to test the model’s long-context capabilities. Moreover, we also reveal that long-context continual pretraining primarily adjusts RoPE embedding to accommodate extended context lengths, which in turn yields only marginal improvements in the model’s true capabilities. To sum up, compared to previous synthetic benchmarks, LongBioBench achieves a better trade-off between mirroring authentic language tasks and maintaining controllability, and is highly interpretable and configurable.  [^1] The code and data are available at https://github.com/Thomasyyj/LongBio-Benchmark.
@@ 1 Introduction
Existing long-context evaluation benchmarks are primarily of two types: natural and synthetic. In terms of natural benchmarks,  Karpinska et al. [3] collects different novels read by annotators and asks them to annotate true/false of the narrative phenomenon.  Li et al. [18] collects the latest documents post year 2022 and crafts questions for them. Though they provide authentic language samples, they are expensive to collect and annotate and susceptible to data contamination. Moreover, their inherent complexity makes them hard to characterize for controlled studies. Thus, the model’s performance on such benchmarks often fails to shed sufficient light on the model’s underlying bottlenecks. In other words, while the data is genuine, it does little to pinpoint precisely how and why a model might struggle, offering limited insights into how its long-context capabilities can be improved. Moreover, since their questions are crafted by human annotators, the task is frozen after annotation and thus unextensible to new evaluation scenarios.
On the contrary, synthetic benchmarks are highly controllable. One can isolate specific aspects they want to test with carefully designed synthetic tasks. Benchmarks like RULER [6] allow practitioners to systematically adjust variables and examine their targeted hypotheses about the model’s potential behaviors. However, existing synthetic benchmarks mostly follow the Needle-In-A-Haystack (NIAH) [19] format, where the context is usually non-coherent because the needle (the information to be recalled) is semantically unrelated to the continuous context. We show in Sec. 4.2 that this makes the benchmark less challenging when the tasks become more difficult, as it potentially presents shortcuts for the model to retrieve the target information. Meanwhile, the complexity of the needle itself is another limitation. As we will show in Sec. 4.1, the numerical needle, which is employed by NIAH [19] and RULER [6] benchmarks, is much easier to retrieve than other types of information. This implicit bias makes them worse proxies of real-world applications, hindering the development of stronger models.
@@
Inspired by Allen-Zhu and Li [20], this paper proposes LongBioBench, employing fictional biographies as a playground to examine existing long-context language models. In our framework, a single configurable biography (the “needle”) is seamlessly embedded within a set of other configurable biographies (the “haystack”), enabling diverse and scalable task design grounded in each bio’s rich factual content. Notably, we show in Fig. 1 that the evaluation scores on our purely synthetic benchmark exhibit a stronger correlation (0.853) with the scores of HELMET [4], which employs real-world tasks, than with other existing purely synthetic dataset RULER [6] (0.559). The task-specific correlations and the analysis are provided in App. G.1.
@@ 2.2 Data Construction
We use the simplest version, referred to as Standard version, as an example to show how a single data point is constructed. An illustration of our proposed context is shown on the left of Fig. 2, and the detailed construction progress is explained in App. B. Specifically, given a task, we first generate needle and haystack biographies using a biography generator. The needle biography is inserted into the haystack to form the context. The corresponding questions and answers are generated alongside the needle biographies, yielding a complete data point. For the biography generator, it typically samples six attributes from predefined attribute pools and fills them into human-written templates to produce coherent biographies for each individual. We also provide flexible configuration options at each step of the construction process to ensure controllability. For instance, we can adjust the distractor density when generating the haystack bios or define the number of required needles when constructing the needle bios. We report the statistics of the dataset in App. C.
The proposed benchmark closely matches the desiderata aforementioned. For seamless context, our generated biographies are ensured to be naturally fluent and maintain the coherence of the entire context. Regarding controllability, the benchmark is highly modular and configurable, allowing us to do isolated ablation studies to probe the model’s behavior. The benchmark is also extensible, as the richly factual nature of the context enables a wide range of downstream tasks, depending on how the embedded knowledge is manipulated (further discussed in the next section). A challenging extended task: In-context Learning is introduced in Sec. 4.5 as an example. Finally, we ensure sound evaluation because the final answer in our benchmark can be easily verified through exact match. We also prevent contamination because all biographies are artificial and can be generated on-the-fly, thus avoiding reliance on any parametric knowledge memorized by the model.
@@ 2.3 Task Description
This subsection will explain how we extend to have all the tasks in LongBioBench. The tasks are split into three categories: understanding, reasoning, and trustworthiness, representing the core capabilities required to solve the tasks. An overview of how each task is extended is shown on the right of Fig. 2. The detailed description for each task is in App. D . Examples are illustrated in Table 2.
@@
Table 2: Task Overview for the LongBioBench. Here {Pn} refers to the name of the n-th person. Acc is the accuracy of the exact match.
@@
Long Context Understanding
Long Context Reasoning
Trustworthiness
3 Main Evaluation Results
3.1 Evaluation setup
We evaluate 15 open-source LCLMs supporting more than 128k context lengths, such as Llama [22], Phi [23], Qwen2.5 [24], Mistral [25]. Details are summerized in App. F. We also include three closed-source models in the GPT family [26]. Each model is evaluated at input lengths: $L\in\{2K,8K,16K,32K,64K,128K\}$ where L is the number of tokens counted with the tokenizer of each model [^2] We did not perform experiments for longer context at scale since most models already perform bad at 128k context length. We provide the results using 256k and 512k context length on three best models in App.G.3. We use zero-shot prompts in all understanding and IDK tasks, and use 2-shot prompting for all reasoning and citation tasks to ensure that the model follows the answer format. We list the prompts for all tasks in App. I. During initial studies, we observed that the model’s performance on our proposed tasks nearly converges  [^3] The accuracy fluctuates within 1% for 32 data points at around 800 data points, so each test set contains 800 data points. We use the vLLM [27] framework for inference with $\timesH8$00 GPUs. We use greedy decoding for all models. Regarding evaluation metrics, we use accuracy with exact match for all understanding and reasoning tasks (all-or-nothing accuracy is applied in the multiple retrieval case). We measure the accuracy of the citations only in the citation task. For the IDK task, the metric is the proportion of questions where the model answers correctly when information is present and refuses to answer when it is absent. We also provide a simple analysis on the hallucination rate in App.G.2.
3.2 Validating all proposed tasks
3.3 Challenges for current long context modeling
This section summarizes six key results observed when evaluating LCLMs on LongBioBench. We first analyze the overall performance across all LCLMs from Fig. 3 (R1 and R2) and look deeper into the performance of different tasks in Fig. 4 (R3-R6). The full results are shown in App. H.
R3: Context length is still the main bottleneck.
R4: Poor correlation between calculation and other reasoning tasks
R5: LCLMs struggle on constrained planning problem
We construct twodiff as an example of a hard task for LCLMs. The answer to this task is not unique, and all bios in the context could serve as the needle. As shown in Fig. 4, even models with strong multihop and arithmetic reasoning abilities—such as Qwen-2.5-14B-1M—struggle with constrained retrieval, failing to perform well even at a 2k context length. Besides, none of the models perform over 30% accuracy at 128k context level, which could be easier with a bigger selection pool. This aligns with the findings from [5] that LCLMs only use a small part of context when doing the reasoning task. This highlights a fundamental limitation: current LCLMs remain far from being able to reason effectively over long contexts.
4 Analysis
4.2 Coherent context is important
4.3 Performance trend during long-context continual pre-training
Setting
Performance saturates at the early stage.
No noticeable improvements on the reasoning abilities are gained through long-ciontext continual pretraining.
Comparing the middle figure with the left one, we observe that performance improves consistently across all retrieval tasks, whereas the reasoning task shows only a slight improvement, with accuracy remaining extremely low at around 10% (note that the random guess accuracy for the ranking task is 50% since it involves ranking two individuals). Interestingly, the calculation task follows a performance trajectory similar to all retrieval tasks and already achieves high accuracy before long-context pretraining. This suggests that Qwen2.5 already possesses the capability to perform calculations over relatively long contexts, and that the main bottleneck on this task lies in retrieval rather than reasoning. Therefore, we categorize the calculation task alongside the retrieval tasks.
The model becomes less trustworthy as the training proceeds
4.4 Distractor density is another bottleneck for long context tasks
To further investigate the factors influencing the performance of LCLMs, we conduct a stress test on Qwen2.5-Instruct-7B-1M by controlling the position of the answer information and the density of distractors as the variables while keeping the context length and task fixed. Detailed analysis is shown in App. E and we summarize the results as follows: (1) We observe a strong negative correlation between distractor density and model performance, suggesting that beyond context length, higher distractor density is a key factor contributing to the difficulty LCLMs face with long-context tasks. (2) We observe the lost-in-the-middle [29] but it is less evident on relatively easier tasks.
4.5 In-context Learning (ICL): An Example for Task Extension
As the foundation capabilities of LLMs wiil keep evolving, constructing more creative and challenging reasoning tasks becoming increasingly essential. In this section, We give a simple example on customizing a more challenging task using LongBioBench.
Inspired by [30], which explores long in-context reasoning capabilities, we add a setting similar to the extreme label task: E.g. “Bio1, Bio2. Question: Which category of university did Charlotte Farley Hall graduate from? Answer: Category 3. … Bio-n, Bio-n+1 Question: Which category of university did Gabriella Jenson Griffin graduate from?”
In this setting, the context comprises multiple in-context demonstrations, each consisting of several biographies and a corresponding QA pair. For each QA pair, we construct a mapping between university names and categorical labels based on their initial letter (e.g., universities starting with “U” are assigned to Category 1, those starting with “A” to Category 3, etc.). Importantly, this mapping is not explicitly provided to the model. Instead, LCLMs are expected to infer the underlying rule by generalizing from the in-context examples. To prevent LCLMs from exploiting memorization or retrieval of the same university name, we ensure that the queried university is unique. However, to support reasoning, we guarantee that at least one university in the demonstrations shares the same initial character, providing a subtle hint for the mapping.
We perform experiments over 10 uniformly distributed categories on the best model Qwen2.5-7b-Instruct-1M and Qwen2.5-14b-Instruct-1M with a cot prompt. The results are shown in Table 3.
Table 3: Results across different ICL context lengths.
ICL 2k 8k 16k 32k 64k 128k
Qwen2.5-14B-Instuct-1M 51.5 25.5 30.5 25.0 16.0 17.0
Qwen2.5-7B-Instuct-1M 20.0 19.5 11.0 11.5 10.5 12.5
Random Guess 10.0 10.0 10.0 10.0 10.0 10.0
We observe several noteworthy findings regarding the performance of LCLMs under varying context lengths. First, as context length increases, models are provided with more demonstrations to learn from; however, we still find substantial performance drops—for instance, from 8k to 16k tokens for Qwen-7B and from 2k to 8k tokens for Qwen-14B—indicating that in-context learning abilities can be negatively affected when handling longer inputs. Second, even state-of-the-art LCLMs continue to struggle with maintaining robust in-context learning over extended contexts. Thirdly, analysis of failure cases reveals a recurring pattern: models often hallucinate category mappings during their reasoning process despite exhibiting strong retrieval capabilities. This behavior highlights a potentially promising optimization direction for improving current LCLMs, particularly those designed with explicit reasoning mechanisms.
5 Conclusion and Limitations
Appendix A Related Work
With the growing interest in long-context language modeling, various benchmarks have emerged. Some focus on real-world tasks across diverse domains, such as document understanding [3, 35, 28], safety [36], and medical question answering [37]. Others try to collect a collection of real-world tasks aimed at a comprehensive evaluation [2, 38, 4]. However, these benchmarks are often costly to construct and suffer from limited interpretability, as it is challenging to control task complexity systematically. On the other hand, many synthetic tasks are proposed due to their low construction cost and high flexibility [6, 39, 31, 5, 7, 40]. Among them, Needle-in-a-Haystack (NIAH) [19] is the most popular setting, where a “needle” is inserted at a long essay (i.e., the haystack) and the model is tasked with retrieving it. There are also variants of NIAH such as RULER [6]. Nonetheless, several studies have raised concerns about the limitations of this approach [5, 8]. In this work, we argue that the semantic irrelevance between the needle and the surrounding context may allow models to exploit cues or shortcuts, ultimately reducing the challenge and bringing bias into evaluations. We note that OpenAI-MRCR [9] also features the blending between needles and haystacks. However, it is less extensible than LongBioBench as the reasoning pattern required to answer is fixed (e.g find the 2nd poem about tapirs), which can be regarded as a specific instantiation of an extension of LongBioBench.
Appendix B Dataset Construction Details
Attribute Sampling
Bio construction
Context Synthesis
Appendix C Data Statistics
Task/Length 2 8 16 32 64 128
standard 14.93 73.77 152.19 308.83 622.18 1248.75
paraphrase 12.88 62.36 128.01 263.86 528.61 1060.80
pronoun 15.07 75.21 156.65 316.81 636.10 1276.81
multi_standard 12.05 70.87 149.28 305.35 617.94 1246.20
calculation 12.84 76.34 161.12 330.70 668.71 1348.31
rank 12.86 75.77 160.62 329.90 667.42 1345.70
multihop 12.06 70.88 149.27 305.82 619.39 1246.22
twodiff 12.85 76.31 161.42 330.84 670.33 1347.75
Task/Length 2 8 16 32 64 128
paraphrase $126.85_{12.68}$ $125.16_{12.15}$ $125.59_{11.60}$ $123.48_{11.45}$ $4_{11.23} 124.1411.2$$4_{11.21}$
Appendix D Task Description
The tasks are split into three categories: understanding, reasoning, and trustworthiness, representing the core capabilities required to solve the tasks. Detailed examples of the benchmark are presented in Table 2.